How much ice at -10C must be added to 0.50 Kg of water at 20 C if it all to melt leaving only water at 0 C?
1
Expert's answer
2011-02-23T06:56:45-0500
Let's construct heat ballance equation:
Cice*m*(T1-T0) + Cfnice * m = Cwater*mwater*(T2-T0)

where Cice is a specific heat of ice, 2.108 J/gK , Cwater is a specific heat of water of 4.187 J/gK, Cfnice - the heat of ice fusion 333.55 J/g, T0 = 0C (melting point of ice), T1 = -10C, T2 = 20C, mwater = 0.5kg=500g.
21.08m + 333.55 m = 41870

m = 118.1 g = 0.118 kg of ice.